1. Understanding the arc elasticy of demand a Aa Suppose that during the past year, the price of a large-screen television rose from $2,000 to $2,600 per television. During the same time period, consumer sales decreased from 700,000 to 650,000 televisions. Calculate the elasticity of demand between these two points by following the steps shown next. After each step, select the appropriate answer Original Quantity New Quantity Average Quantity Change in Quantity Percentage Change in Quantity Original Price New Price Average Price Change in Price Percentage Change in Price Step 1: Select the correct answers for original quantity, new quantity, original price, and new price Step 2: Calculate the average quantity by adding the original quantity and new quantity, and then divide by two. Do the same for the average price. Step 3: Calculate the change in quantity by subtracting the original quantity from the new quantity. Do the same for the change in price. Step 4: Calculate the percentage change in quantity demanded by dividing the change in quantity by the average quantity. Do the same to calculate the percentage change in price. Step 5: Calculate the price elasticity of demand by dividing the percentage change in quantity demanded by the percentage change in price. The elasticity of demand for large-screen televisions is about O-3.53 O-4.20 O-0.28 O-0.24

Hello - There are 11 parts to This question From left to right the multiple choice options are listed below:

1. (700,000, 2,000, 2,600, 650,000)

2. (700,000, 2,000, 2,600, 650,000)

3. (2,300, 1,3500,000, 4,600, 675,000)

4. (-50,000, -600)

5. (-7.1%, -7.4%, -30%)

6. ($650,000, $700,000, $2,000, $2,600)

7. ($650,000, $700,000, $2,000, $2,600)

8. ($4,600, $1,350,000, $2,300, $675,000)

9. ($600, $50,000)

10. (26.1%, 30%, 7.1%)

11. (-3.53, -4.20, -.28, -.24)

Thank you for your help!
Answer #1

1) Original Quantity = 700,000

2) New Quantity = 650,000

3) Average Quantity = (700,000+650,000)/2 = 675,000

4) Change in Quantity = (650,000-700,000) = -50,000

5)% change in Quantity = (-50,000/700,000)*100 = -7.1%

6) Original price = $2000

7) New price = $2600

8) Average price = (2000+2600)/2 = $2300

9) Change in price = 2600-2000 = $600

10) % change in price = (600/2000)*100 = 30%

11) Elasticity of demand = % change in quantity/ % change in price = -7.1/30 = -0.24